# Is Our Chatbot Telling Lies? Assessing Correctness of an LLM-based Dutch Support Chatbot

### v1

@@ IV-C1 Message Types
Since the type of message impacts the nature of the response, it is essential to identify the various message types. This way, we can examine how truthfulness can be assessed for each type. As far as we are aware, there is no existing classification of user message types, except those related to financial inquiries [19]. In this study, the authors utilize message type to tailor an LLM prompt to the financial objective of the question. They opted to generate question labels with particular intentions. On the other hand, our taxonomy is crafted to be more general, not delving into the precise nature of financial questions but rather focusing on the type of information sought. This enables us to be adaptable to previously unseen questions.
We analyzed a random sub-sample of approximately 300 messages sent to the chatbot to create the classification. These messages are message-answer pairs that were not rated, ensuring they are not subject to selection bias by the annotators [49]. Based on this analysis, we identified seven types of user messages sent by the support team. These types are derived from the decision tree and those observed in the random subset. They are: (From now on, the underlined names will refer to these types.)
@@
The message types for both the analysis set and test set are labeled by the first author. The author’s labeling is cross-validated by three AI developers of AFAS, who each annotate a random subset of messages, covering 77% of the complete analysis set. The inter-annotator agreement is computed using Cohen’s kappa [50], with values below 0 indicating disagreement, above 0 agreement, and 1 perfect agreement. Kappa has been used in similar research [51, 41, 35, 18], with resulting values often ranging between 0.3 and 0.5 [41], and has been employed to evaluate the usefulness of error taxonomies for chatbots reaching a kappa of 0.44 [51]. Our message type taxonomy achieves a Cohen’s kappa of 0.65, which is considered moderate [35]. The author and developers have a high agreement of 0.81 for Binary and Instruction message types, with 91% of messages consistently labeled. Given the overlap between the two types, using them together for automated scoring makes sense.
IV-C2 Heuristics identified
@@
Table II shows the heuristics observed for each message type. This further justifies our choice to study Binary and Instruction types, given a substantial overlap of heuristics to measure truthfulness.
@@
TABLE II: Overview of heuristics for each type of message. From syntactic to semantic heuristics.
@@
Figure 3: Total number of message-answer pairs per type of mistake. On the x-axis, the type of message is shown, and the y-axis shows the total number of messages classified per each type.
V From Human to Automated Assessment
@@
V-A Identify Type of User Message
Due to the limited number of messages and the presence of 7 different message types, only a few samples per class are available. E.g. only 5 for type Error, see Figure 3. While it is common to train a machine learning model for classification [52], our sparse dataset is not large enough for both training and testing. Consequently, we attempt to develop a rule-based algorithm, based on human analysis, to identify the types of messages.
@@ V-A Identify Type of User Message
In natural language processing it is common to employ pre-processing before the actual text processing, such as classification. Examples of these techniques include stopword removal [53, 54, 55], lemmatization [56], punctuation removal [53, 55, 57] and lowercasing [53, 54, 55, 57]. These methods help overcome minor variations between identical words or sentences, reducing noise in the data. For the message type prediction, the text is lowercased as the lists with patterns are lowercased and not case-sensitive.
For a message, our approach scans for words or patterns from the first list. If a match is found, a message is assigned to the corresponding message type. If no match is found, the system checks for a match in the next list. This is continued till a match is found and the type unspecified is assigned otherwise. Therefore, if a question contains words from multiple lists, it is assigned the type of the first list with which it shares a word. The order of the lists is determined based on their performance on the analysis set, focusing especially on the accuracy of classifying Binary and Instruction types. This is because the main interest is in accurately classifying these two types, as only those types will be scored. This leads to the following sequence of lists; for complete lists, please refer to the replication package [1]:
@@
V-B Feature Selection
With truthfulness defined and the heuristics for a true answer established, the next step is transforming heuristics into automated features. Note that LLM not having the knowledge about the product is a key factor in this research. As the external trained model has no knowledge about the company, all info mentioned should be present in the context, otherwise it is hallucinating [19]. Therefore, we add verifying whether the content is contained within the context to the heuristics.
To translate the heuristics to automated features, we use literature to find features to automate this. To find relevant papers, the words from the following non-exhaustive list are combined: metric, correctness, score, evaluate, chatbot, conversational bot, conversational agent, hallucination, characteristics, wrong, correct, NLP, education, grading, measure, automated. Using the word combinations, papers are found utilizing various scientific databanks including ResearchGate [^3] https://www.researchgate.net/, Google Scholar [^4] https://scholar.google.com/, and IEEE Xplore [^5] https://ieeexplore.ieee.org/Xplore/home.jsp. Additionally, relevant papers are used for snowballing, both backward and forward. About 70 relevant papers are identified, which either define a non-automated metric for chatbots or an automatic metric for general text grading. We identified two relevant fields with related features: automated metrics for chatbots and automated grading of student answers. A full list is constructed with potential relevant features, see the replication package [1] for all identified features.
@@ V-B Feature Selection
Each of the implemented features is assessed for Spearman correlation [58] with human evaluation. If a feature has a positive correlation and a p-value lower than 0.10, it is selected for the final selection. While a correlation of 0.10 is not statistically significant it is deemed sufficient, as the actual selection will take place during the final selection. If a feature doesn’t meet the correlation criteria but has demonstrated effectiveness in relevant literature, it is tested if it distinguishes between true and untrue answers. If it can differentiate at least some of these answers, it is selected for the final selection. Additionally, in the final selection, features that perform the exact same function as the heuristic they are intended to replace, such as verifying the presence of a word, are included.
After this raw selection of most promising features, the final selection is done using an ablation study [59, 60]. We assess whether removing the feature impacts the significant correlation between the combined features and the human evaluation for the analysis set. If the removal decreases this correlation, we keep the feature. This hierarchy of filtering steps enable an initial rough selection with limited confidence, as features are ultimately chosen only if they pass the strictest selection criterion, which is the final step.
@@
$\rho$ Company-Specific Terms - This feature is developed based on the General Answer heuristic. The assumption is that an answer should not include general terms when describing a solution. The 10,000 most frequent used words in AFAS its help documentation are checked against all words in the Dutch Wikibooks dataset [61]. Words not found in this dataset are considered company-specific, and answers with such words are deemed more truthful.
Components Defined - As per the Unspecified Component heuristic, an answer should only include existing components. Which is influenced by the feature introduced by Roychowdhury [19], where they verify the precise financial numbers within the context. In the help documentation, components are defined structurally, and the LLM preserves this structure in its responses, even if mentioning non-existing components. Therefore, components can be extracted using REGEX. If an answer includes components not defined in the context, it is less likely to be true.
$\rho$ [Uncaptioned image]Complex Answer - Following the General Answer heuristic, we check for the existence of words indicating a complex text, as introduced by Kumar et al. [16]. While no such lists are available in scientific literature for Dutch, we created four lists based on signal words from Genootschap Onze Taal666(Society Our Language): https://onzetaal.nl/taalloket/signaalwoorden-lijst (29-05-2024) and Boom NT2777https://www.nt2.nl/documenten/luisteren_op_b2/overzicht_van_ signaalwoorden.pdf (29-05-2024), each corresponding to different types of complexity: perspective, comparison, examples, and reasoning. Full lists can be found in the replication package [1]. The presence of words from various lists increases the likelihood that the answer is true.
Prompt Overlap - Roychowdhury et al. introduced the Prompt Uniqueness feature, which indicates that answers are less likely to be good if they repeat parts of the question [19]. In contrast, Kumar et al. proposed Prompt Overlap, suggesting that some overlap is expected and, when present, indicates a better answer [16].
@@
$\rho$ [Uncaptioned image] [Uncaptioned image]HAL - In order to address the Off-Context heuristic, the HAL technique from the study by Lund et al. [62] is implemented. This feature measures how often pairs of words appear together within a sliding window of varying sizes, giving higher scores to pairs that frequently occur next to each other. The intuition is that if a word or setting frequently appears together in the answer, it should also be close together in the context. If not, it is less likely a true answer.
@@
V-C Features to Score
To go from these features to a score, all features are normalized between 0 and 1. Following the method of Roychowdhury et al. [19], we sum the features together. We normalize the sum between 1 and 5, adhering to the rating of human annotators.
@@ V-C Features to Score
Figure 4: A visual explanation of how the score is derived.
VI Evaluation
@@ VI-A Method of Evaluation
Evaluation on the scoring is done with the Dutch text and English text. For the Dutch text, tests showed that the features worked best when lemmatized and lowercased. In section V-B the [Uncaptioned image]symbol shows for which features English text is used in the English scoring version. These features utilize externally trained packages, which work with English text as well. The other features still use Dutch text because they rely on custom words and regex fine-tuned for Dutch. The English texts show the best result when lemmatized, lowercased and stripped from stopwords.
The evaluation of the resulting scores is conducted across three gradations: overall performance, variations in prediction, and situations where the approach is ineffective. For the overall performance, the Spearman correlation between the prediction and the human evaluation is used. This is commonly used in similar research [33, 29, 63]. With significance levels of 0.05 [32, 35] and 0.01 [64, 29]. Secondly, we assess the deviations using an error margin. As the rating of the human annotators is discrete, and our metric continuous. An error margin of 1 is used, as it reflects whether a positive rated message ($>3$) is rated above three by our metric, and vice versa. Finally, the scenarios for which our metric does not work are determined by manually assessing mistakes made.
VI-B Results
VI-B1 Message Type Prediction
Table III shows the evaluation of the type prediction for both, analysis and test set. The overall performance is shown, along with the predictions for the types Binary and Instruction. All patterns utilized in the prediction are derived solely from the analysis set. This may result in many messages in the test set not matching any pattern; however, only 18% of the test set is labeled as unspecified, suggesting that at most 18% of the messages contain unseen patterns. Among the messages labeled as Binary and Instruction, only six messages do not match any pattern.
Bear in mind that the prediction is solely based on rule-based checks. After all, we achieve an F1 score of 0.77 on the test set for type Binary and Instruction, and an F1 score of 0.80 for Instruction type. Table III displays that 82% of the messages that will be scored are genuinely of type Binary or Instruction. Additionally, around 72% of messages identified as Binary and Instruction are accurately predicted as such, and consequently scored, as indicated in Table III.
@@ VI-B1 Message Type Prediction
TABLE III: Automated message detection performance on training and testing datasets: all for overall classification, Binary and Instruction for labeling respective types, and is Binary or Instruction for distinguishing either type or neither.
VI-B2 Overall
The overall performance indicates that the Dutch features have been tailored and optimized for the analysis set, as they exhibit better performance on this set compared to the English features and the Test set, as shown in Table IV. However, this is not the case for the translated version, where the test set outperforms the analysis set. All of the features are fine-tuned on the Dutch data, therefore it is possibly overfitted. For the test set, it concludes that the translation has a positive effect on the prediction. This indicates that the external packages used work better with English than with Dutch text.
To contextualize the correlation of our metric, we refer to the most comparable metric in terms of definition discovered in the literature, developed by Mehri et al. [20]. They evaluate answers based on correctness and achieve a correlation of 0.13. Notably, the Dutch and English versions exceed this result, with correlations of 0.28 and 0.37, respectively.
@@ VI-B2 Overall
TABLE IV: Spearman Correlation between Human truthfulness evaluation and automated Score results. * for p$<$0.05, ** for p$<$0.01.
VI-B3 Deviations
To assess the accuracy and deviation, see Table V. It performs especially well in rating the high and low rated messages, not the neutral. Additionally, Table V shows that over half of those rated with 1 star are containing non-existing components. Notably, 60% of neutral responses include a non-existent component. This may imply that the answer is true and comprehensible. However, due to the incorrect terminology used, the answer cannot be considered entirely true. Table V shows a threshold of 3, which indicates neutrality. Messages with a score above 3 should be considered true, and those below 3 should be considered untrue. As demonstrated, 67% of the 1-star rated messages have a score below three, while only 21% of the 5-star rated messages do. Therefore, by not sending any messages with a score lower than 3 to the user, 64% of the messages do not need to be judged by hand by the support team, and only 21% of the 5-star rated messages will be discarded.

### v2

@@ 4.4.1 Message Types
Since the type of message impacts the nature of the response, it is essential to identify the various message types. This way, we can examine how truthfulness can be assessed for each type. As far as we are aware, there is no existing classification of user message types, except those related to financial inquiries (Roychowdhury et al., 2023) or chatbots for general use (Shah et al., 2024; Ji et al., 2023b). In this study, the authors utilize message type to tailor an LLM prompt to the financial objective of the question. They opted to generate question labels with particular intentions. On the other hand, our taxonomy is crafted to be more general for support bots in any field of business, not delving into the precise nature of financial questions but rather focusing on the type of information sought. This enables us to be adaptable to numerous companies and to previously unseen or unexpected questions.
We analyzed a random sub-sample of approximately 300 messages sent to the chatbot to create the classification. These messages are message-answer pairs that were not rated, ensuring they are not subject to selection bias by the annotators (Gupta et al., 2024). Based on this analysis, we identified seven types of user messages sent by the support team. These types are derived from the decision tree and those observed in the random subset. They are: (From now on, the underlined names will refer to these types.)
@@
The message types for both the analysis set and test set are labeled by the first author. The author’s labeling is cross-validated by three AI developers of AFAS, who each annotate a random subset of messages, covering 77% of the complete analysis set. The inter-annotator agreement is computed using Cohen’s kappa (McHugh, 2012), with values below 0 indicating disagreement, above 0 agreement, and 1 perfect agreement. Kappa has been used in similar research (Higashinaka et al., 2021; van der Lee et al., 2019; Lowe et al., 2017; Merdivan et al., 2020), with resulting values often ranging between 0.3 and 0.5 (van der Lee et al., 2019), and has been employed to evaluate the usefulness of error taxonomies for chatbots reaching a kappa of 0.44 (Higashinaka et al., 2021). Our message type taxonomy achieves a Cohen’s kappa of 0.65, which is considered moderate (Lowe et al., 2017). The author and developers have a high agreement of 0.81 for Binary and Instruction message types, with 91% of messages consistently labeled. Given the overlap between the two types, using them together for automated scoring makes sense.
4.4.2 Heuristics identified
@@
Table 2 shows the heuristics observed for each message type. This further justifies our choice to study Binary and Instruction types, given a substantial overlap of heuristics to measure truthfulness.
@@
Table 2: Overview of heuristics for each type of message. Ranging from syntactic to semantic heuristics. x-axis showing the type of message, y-axis the heuristic
@@
Figure 3: Total message-answer pairs by mistake type. The x-axis shows the message type, and the y-axis shows the number of messages per type
5 From Human to Automated Assessment
@@
5.1 Identify Type of User Message
Due to the limited number of messages and the presence of 7 different message types, only a few samples per class are available. E.g. only 5 for type Error, see Figure 3. While it is common to train a machine learning model for classification (Sen et al., 2019), our sparse dataset is not large enough for both training and testing. Consequently, we attempt to develop a rule-based algorithm, based on human analysis, to identify the types of messages.
@@ 5.1 Identify Type of User Message
In natural language processing it is common to employ pre-processing before the actual text processing, such as classification. Examples of these techniques include stopword removal (Obie et al., 2023; Carreno and Winbladh, 2013; Zhan et al., 2022), lemmatization (Phong et al., 2015), punctuation removal (Obie et al., 2023; Zhan et al., 2022; Tahvili et al., 2020) and lowercasing (Obie et al., 2023; Carreno and Winbladh, 2013; Zhan et al., 2022; Tahvili et al., 2020). These methods help overcome minor variations between identical words or sentences, reducing noise in the data. For the message type prediction, the text is lowercased as the lists with patterns are lowercased and not case-sensitive.
For a message, our approach scans for words or patterns from the first list. If a match is found, a message is assigned to the corresponding message type. If no match is found, the system checks for a match in the next list. This is continued till a match is found and the type unspecified is assigned otherwise. Therefore, if a question contains words from multiple lists, it is assigned the type of the first list with which it shares a word. The order of the lists is determined based on their performance on the analysis set, focusing especially on the accuracy of classifying Binary and Instruction types. This is because the main interest is in accurately classifying these two types, as only those types will be scored. This leads to the following sequence of lists; for complete lists, please refer to the replication package (Lassche et al., 2024):
@@
5.2 Feature Selection
With truthfulness defined and the heuristics for a true answer established, the next step is transforming heuristics into automated features. Note that LLM not having the knowledge about the product is a key factor in this research. As the external trained model has no knowledge about the company, all info mentioned should be present in the context, otherwise it is hallucinating (Roychowdhury et al., 2023). Therefore, we add verifying whether the content is contained within the context to the heuristics. At the beginning of our study, we made the decision to avoid using LLMs for answer evaluation. This decision was based on existing evaluations with LLMs, which have proven to be expensive and time-consuming (Lin and Chen, 2023). Our goal is to provide real-time scoring so that users do not have to wait for an additional model call. Additionally, LLMs are inconsistent in scoring and error-prone, often producing different outcomes depending on the prompt and model used (Lin and Chen, 2023). We also aim for consistent scoring for each question-answer pair to ensure reliable comparisons.
To translate the heuristics to automated features, we use literature to find features to automate this. To find relevant papers, the words from the following non-exhaustive list are combined: metric, correctness, score, evaluate, chatbot, conversational bot, conversational agent, hallucination, characteristics, wrong, correct, NLP, education, grading, measure, automated. Using the word combinations, papers are found utilizing various scientific databanks including ResearchGate [^3] https://www.researchgate.net/, Google Scholar [^4] https://scholar.google.com/, and IEEE Xplore [^5] https://ieeexplore.ieee.org/Xplore/home.jsp. Additionally, relevant papers are used for snowballing, both backward and forward. About 70 relevant papers are identified, which either define a non-automated metric for chatbots or an automatic metric for general text grading. We identified two relevant fields with related features: automated metrics for chatbots and automated grading of student answers. A full list is constructed with potential relevant features, see the replication package (Lassche et al., 2024) for all identified features.
@@ 5.2 Feature Selection
Each of the implemented features is assessed for Spearman correlation (Spearman, 1905) with human evaluation. If a feature has a positive correlation and a p-value lower than 0.10, it is selected for the final selection. While a correlation of 0.10 is not statistically significant it is deemed sufficient, as the actual selection will take place during the final selection. If a feature doesn’t meet the correlation criteria but has demonstrated effectiveness in relevant literature, it is tested if it distinguishes between true and untrue answers. If it can differentiate at least some of these answers, it is selected for the final selection. Additionally, in the final selection, features that perform the exact same function as the heuristic they are intended to replace, such as verifying the presence of a word, are included.
After this raw selection of most promising features, the final selection is done using an ablation study (Sheikholeslami, 2019; Meyes et al., 2019). We assess whether removing the feature impacts the significant correlation between the combined features and the human evaluation for the analysis set. If the removal decreases this correlation, we keep the feature. This hierarchy of filtering steps enable an initial rough selection with limited confidence, as features are ultimately chosen only if they pass the strictest selection criterion, which is the final step.
@@
$\rho$ Company-Specific Terms - This feature is developed based on the General Answer heuristic. The assumption is that an answer should not include general terms when describing a solution. The 10,000 most frequent used words in AFAS its help documentation are checked against all words in the Dutch Wikibooks dataset (Dave, 2021). Words not found in this dataset are considered company-specific, and answers with such words are deemed more truthful.
Components Defined - As per the Unspecified Component heuristic, an answer should only include existing components. Which is influenced by the feature introduced by Roychowdhury et al. (2023), where they verify the precise financial numbers within the context. In the help documentation, components are defined structurally, and the LLM preserves this structure in its responses, even if mentioning non-existing components. Therefore, components can be extracted using REGEX. If an answer includes components not defined in the context, it is less likely to be true.
$\rho$ [Uncaptioned image]Complex Answer - Following the General Answer heuristic, we check for the existence of words indicating a complex text, as introduced by Kumar et al. (2019). While no such lists are available in scientific literature for Dutch, we created four lists based on signal words from Genootschap Onze Taal666(Society Our Language): https://onzetaal.nl/taalloket/signaalwoorden-lijst (29-05-2024) and Boom NT2777https://www.nt2.nl/documenten/luisteren_op_b2/overzicht_van_ signaalwoorden.pdf (29-05-2024), each corresponding to different types of complexity: perspective, comparison, examples, and reasoning. Full lists can be found in the replication package (Lassche et al., 2024). The presence of words from various lists increases the likelihood that the answer is true.
Prompt Overlap - Roychowdhury et al. (2023) introduced the Prompt Uniqueness feature, which indicates that answers are less likely to be good if they repeat parts of the question. In contrast, Kumar et al. (2019) proposed Prompt Overlap, suggesting that some overlap is expected and, when present, indicates a better answer.
@@
$\rho$ [Uncaptioned image] [Uncaptioned image]HAL - In order to address the Off-Context heuristic, the HAL technique from the study by Lund and Burgess (1996) is implemented. This feature measures how often pairs of words appear together within a sliding window of varying sizes, giving higher scores to pairs that frequently occur next to each other. The intuition is that if a word or setting frequently appears together in the answer, it should also be close together in the context. If not, it is less likely a true answer.
@@
5.3 Features to Score
To go from these features to a score, all features are normalized between 0 and 1. Following the method of Roychowdhury et al. (2023), we sum the features together. We normalize the sum between 1 and 5, adhering to the rating of human annotators.
@@ 5.3 Features to Score
Figure 4: Illustration of the scoring process, where scores range from 1 to 5. A score of 1 is assigned if any component is missing, and 5 if a guide in the answer matches one in the context. Intermediate scores are obtained by summing the outputs of the other metrics
6 Evaluation
@@ 6.1 Method of Evaluation
Evaluation on the scoring is done with the Dutch text and English text. For the Dutch text, tests showed that the features worked best when lemmatized and lowercased. In section 5.2 the [Uncaptioned image]symbol shows for which features English text is used in the English scoring version. These features utilize externally trained packages, which work with English text as well. The other features still use Dutch text because they rely on custom words and regex fine-tuned for Dutch. The English texts show the best result when lemmatized, lowercased and stripped from stopwords.
The evaluation of the resulting scores is conducted across three gradations: overall performance, variations in prediction, and situations where the approach is ineffective. For the overall performance, the Spearman correlation between the prediction and the human evaluation is used. This is commonly used in similar research (Pang et al., 2020; Tao et al., 2018; Haque et al., 2022). With significance levels of 0.05 (Huang et al., 2020; Lowe et al., 2017) and 0.01 (Guan and Huang, 2020; Tao et al., 2018). Secondly, we assess the deviations using an error margin. As the rating of the human annotators is discrete, and our metric continuous. An error margin of 1 is used, as it reflects whether a positive rated message (>3>3) is rated above three by our metric, and vice versa. Finally, the scenarios for which our metric does not work are determined by manually assessing mistakes made.
6.2 Results
6.2.1 Message Type Prediction
Table 3 shows the evaluation of the type prediction for both, analysis and test set. The overall performance is shown, along with the predictions for the types Binary and Instruction. All patterns utilized in the prediction are derived solely from the analysis set. This may result in many messages in the test set not matching any pattern; however, only 18% of the test set is labeled as unspecified, suggesting that at most 18% of the messages contain unseen patterns. Among the messages labeled as Binary and Instruction, only six messages do not match any pattern.
Bear in mind that the prediction is solely based on rule-based checks. After all, we achieve an F1 score of 0.77 on the test set for type Binary and Instruction, and an F1 score of 0.80 for Instruction type. Table 3 displays that 82% of the messages that will be scored are genuinely of type Binary or Instruction. Additionally, around 72% of messages identified as Binary and Instruction are accurately predicted as such, and consequently scored, as indicated in Table 3.
@@ 6.2.1 Message Type Prediction
Table 3: Automated message detection performance on training and testing datasets: all for overall classification, Binary and Instruction for labeling respective types, and is Binary or Instruction for distinguishing either type or neither
6.2.2 Overall
The overall performance indicates that the Dutch features have been tailored and optimized for the analysis set, as they exhibit better performance on this set compared to the English features and the Test set, as shown in Table 4. However, this is not the case for the translated version, where the test set outperforms the analysis set. All of the features are fine-tuned on the Dutch data, therefore it is possibly overfitted. For the test set, it concludes that the translation has a positive effect on the prediction. This indicates that the external packages used work better with English than with Dutch text.
To contextualize the correlation of our metric, we refer to the most comparable metric in terms of definition discovered in the literature, developed by Mehri and Eskenazi (2020). They evaluate answers based on correctness and achieve a correlation of 0.13. Notably, the Dutch and English versions exceed this result, with correlations of 0.28 and 0.37, respectively.
@@ 6.2.2 Overall
Table 4: Spearman Correlation between Human truthfulness evaluation and automated Score results. * for p<<0.05, ** for p<<0.01
6.2.3 Deviations
To assess the accuracy and deviation, see Table 5. It performs especially well in rating the high and low rated messages, not the neutral. Additionally, Table 5 shows that over half of those rated with 1 star are containing non-existing components. Notably, 60% of neutral responses include a non-existent component. This may imply that the answer is true and comprehensible. However, due to the incorrect terminology used, the answer cannot be considered entirely true. Table 5 shows a threshold of 3, which indicates neutrality. Messages with a score above 3 should be considered true, and those below 3 should be considered untrue. As demonstrated, 67% of the 1-star rated messages have a score below three, while only 21% of the 5-star rated messages do. Therefore, by not sending any messages with a score lower than 3 to the user, 64% of the messages do not need to be judged by hand by the support team, and only 21% of the 5-star rated messages will be discarded.